OpenStudy (anonymous):

What is F=9/5C+32, Solve for C

OpenStudy (campbell_st):

ok... you need to make C the subject... so subtract 32 from both sides... what do you get?

OpenStudy (anonymous):

\[F-32=\frac{9 }{ 5 }C\]

OpenStudy (campbell_st):

excellent mutiply both sides by 5 to eliminate the fraction then the step after that is to divide both sides by 9

OpenStudy (anonymous):

so it's \[\left( \frac{ 5 }{ 9 } \right)\left( F-32 \right)=C\]

OpenStudy (campbell_st):

thats correct... well done..

OpenStudy (anonymous):

Thanks you!